OpenStudy (anonymous):

Calculate the number of subsets and the number of proper subsets for the set below. M={x|x is an integer between 1 and 7}

jimthompson5910 (jim_thompson5910):

There are 7 numbers in the set M, so there are 2^7 = 128 subsets and 128 - 1 = 127 proper subsets.

OpenStudy (anonymous):

yeah thats what i got but i am on mathxl and its saying its not right

jimthompson5910 (jim_thompson5910):

hmm interesting..maybe they're excluding 1 and 7. So there would be 5 numbers in set M which would mean 2^5 = 32 subsets and 32-1 = 31 proper subsets.

jimthompson5910 (jim_thompson5910):

either that, or there's a typo somewhere

OpenStudy (anonymous):

thats what the answer was. the second one. what would it be if it was 0 and 6

OpenStudy (anonymous):

it would be the same answer right

jimthompson5910 (jim_thompson5910):

There are 5 numbers from 0 to 6 (excluding 0 and 6) since they are: 1, 2, 3, 4, 5 So there are 2^5 = 32 subsets and 32-1 = 31 proper subsets.

jimthompson5910 (jim_thompson5910):

yes exactly

OpenStudy (anonymous):

okay thank you i understand it now. they just give horrible examples in math books

jimthompson5910 (jim_thompson5910):

i see, sounds like a lot of text books i've read lol